OpenStudy (anonymous):

Rationalize the denominator and simplify (y)/(square root (y) - 5 square root (x)

OpenStudy (ash2326):

\[y/(\sqrt(y)-5\sqrt(x))\] multiply numerator and denominator by \[\sqrt(y)+5\sqrt(x)\]

OpenStudy (lalaly):

\[\frac{y}{\sqrt y - 5\sqrt x} \times \frac{\sqrt y+ 5\sqrt x}{\sqrt y+ 5\sqrt x}\]

OpenStudy (ash2326):

so we get \[y(\sqrt y +5 \sqrt x)/(y-25x)\]